WorldClass Statistics · Appendix: Exercise solutions

Solutions: Probability

9.79(a) False. These are independent trials. (b) False. There are red face cards. (c) True. A card cannot be both a face card and an ace.

9.81(a) 10 tosses. Fewer tosses mean more variability in the sample fraction of heads, meaning there’s a better chance of getting at least 60% heads. (b) 100 tosses. More flips means the observed proportion of heads would often be closer to the average, 0.50, and therefore also above 0.40. (c) 100 tosses. With more flips, the observed proportion of heads would often be closer to the average, 0.50. (d) 10 tosses. Fewer flips would increase variability in the fraction of tosses that are heads.

9.83(a) 0.5100.5^{10} = 0.00098. (b) 0.5100.5^{10} = 0.00098. (c) PP(at least one tails) = 1−P1 - P(no tails) = 1−(0.510)≈1−0.001=0.9991 - (0.5^{10}) \approx 1 - 0.001 = 0.999.

9.85(a) No, there are voters who are both independent and swing voters.
(b)

(c) Each Independent voter is either a swing voter or not. Since 35% of voters are Independents and 11% are both Independent and swing voters, the other 24% must not be swing voters. (d) 0.47. (e) 0.53. (f) P(Independent) ×\times P(swing) = 0.35×0.23=0.080.35\times0.23 = 0.08, which does not equal P(Independent and swing) = 0.11, so the events are dependent.

9.87(a) If the class is not graded on a curve, they are independent. If graded on a curve, then neither independent nor disjoint – unless the instructor will only give one A, which is a situation we will ignore in parts (b) and (c). (b) They are probably not independent: if you study together, your study habits would be related, which suggests your course performances are also related. (c) No. See the answer to part (a) when the course is not graded on a curve. More generally: if two things are unrelated (independent), then one occurring does not preclude the other from occurring.

9.89(a) 0.16+0.09=0.250.16 + 0.09 = 0.25. (b) 0.17+0.09=0.260.17 + 0.09 = 0.26. (c) Assuming that the education level of the husband and wife are independent: 0.25×0.26=0.0650.25 \times 0.26 = 0.065. You might also notice we actually made a second assumption: that the decision to get married is unrelated to education level. (d) The husband/wife independence assumption is probably not reasonable, because people often marry another person with a comparable level of education. We will leave it to you to think about whether the second assumption noted in part (c) is reasonable.

9.91(a) No, but we could if A and B are independent. (b-i) 0.21. (b-ii) 0.79. (b-iii) 0.3. (c) No, because 0.1 ≠\ne 0.21, where 0.21 was the value computed under independence from part (a). (d) 0.143.

9.93(a) No, 0.18 of respondents fall into this combination. (b) 0.60+0.20−0.18=0.620.60 + 0.20 - 0.18 = 0.62. (c) 0.18/0.20=0.90.18 / 0.20 = 0.9. (d) 0.11/0.33≈0.330.11 / 0.33 \approx 0.33. (e) No, otherwise the answers to (c) and (d) would be the same. (f) 0.06/0.34≈0.180.06 / 0.34 \approx 0.18.

9.95(a) No. There are 6 females who like Five Guys Burgers. (b) 162/248=0.65162 / 248 = 0.65. (c) 181/252=0.72181 / 252 = 0.72. (d) Under the assumption of a dating choices being independent of hamburger preference, which on the surface seems reasonable: 0.65×0.72=0.4680.65 \times 0.72 = 0.468. (e) (252+6−1)/500=0.514(252 + 6 - 1)/500 = 0.514.

9.97(a)
A tree diagram with a primary branch "Can construct box plots?" and a secondary branch "Passed?". The primary "Can construct box plots" branching has two possibilities of "Yes" with probability 0.8 and "No" with probability 0.2. Each of these branches has two secondary branches. The "Yes" primary branch breaks into branches for "Yes" (for Passed) that has a conditional probability of 0.86 with a Yes-and-Yes final probability of 0.688, and a "No" secondary branch with a conditional probability of 0.14 with a Yes-and-No final probability of 0.112. The "No" primary branch from "Can construct box plots" has a branch of "Yes" that has a conditional probability of 0.65 with a No-and-Yes final probability of 0.13, and a "No" secondary branch with a conditional probability of 0.35 with a No-and-No final probability of 0.07. (b) 0.84

9.99. Even when a patient tests positive for lupus, there is only a 7.14% chance that he actually has lupus. House may be right.

9.101(a) 0.3. (b) 0.3. (c) 0.3. (d) 0.3×0.3=0.090.3\times0.3=0.09. (e) Yes, the population that is being sampled from is identical in each draw.

9.103(a) 2/9≈0.222 / 9 \approx 0.22. (b) 3/9≈0.333 / 9 \approx 0.33. (c) 310×29≈0.067\frac{3}{10} \times \frac{2}{9} \approx 0.067. (d) No, e.g. in this exercise, removing one chip meaningfully changes the probability of what might be drawn next.

9.105P(1P(^1leggings, 2^2jeans, 3^3jeans)=524×723×622=0.0173) = \frac{5}{24} \times \frac{7}{23} \times \frac{6}{22} = 0.0173. However, the person with leggings could have come 2nd or 3rd, and these each have this same probability, so 3×0.0173=0.05193 \times 0.0173 = 0.0519.

9.107(a) 13. (b) No, these 27 students are not a random sample from the university’s student population. For example, it might be argued that the proportion of smokers among students who go to the gym at 9 am on a Saturday morning would be lower than the proportion of smokers in the university as a whole.

9.109(a) E(X) = 3.59. SD(X) = 9.64. (b) E(X) = -1.41. SD(X) = 9.64. (c) No, the expected net profit is negative, so on average you expect to lose money.

9.111% increase in value.

9.113E = -0.0526. SD = 0.9986.

9.115Approximate answers are OK.
(a) (29+32)/144=0.42(29+32)/144 = 0.42. (b) 21/144=0.1521/144 = 0.15. (c) (26+12+15)/144=0.37(26+12+15)/144 = 0.37.

9.117(a) Invalid. Sum is greater than 1. (b) Valid. Probabilities are between 0 and 1, and they sum to 1. In this class, every student gets a C. (c) Invalid. Sum is less than 1. (d) Invalid. There is a negative probability. (e) Valid. Probabilities are between 0 and 1, and they sum to 1. (f) Invalid. There is a negative probability.

9.119.

9.121(a) E = $3.90. SD = $0.34.
(b) E = $27.30. SD = $0.89.

9.123Var(X1+X22)Var\left(\frac{X_1 + X_2}{2}\right)
=Var(X12+X22)= Var\left(\frac{X_1}{2} + \frac{X_2}{2}\right)
=Var(X1)22+Var(X2)22= \frac{Var(X_1)}{2^2} + \frac{Var(X_2)}{2^2}
=σ24+σ24= \frac{\sigma^2}{4} + \frac{\sigma^2}{4}
=σ2/2= \sigma^2 / 2
9.125Var(X1+X2+⋯+Xnn)Var\left(\frac{X_1 + X_2 + \dots + X_n}{n}\right)
=Var(X1n+X2n+⋯+Xnn)= Var\left(\frac{X_1}{n} + \frac{X_2}{n} + \dots + \frac{X_n}{n}\right)
=Var(X1)n2+Var(X2)n2+⋯+Var(Xn)n2= \frac{Var(X_1)}{n^2} + \frac{Var(X_2)}{n^2} + \dots + \frac{Var(X_n)}{n^2}
=σ2n2+σ2n2+⋯+σ2n2= \frac{\sigma^2}{n^2} + \frac{\sigma^2}{n^2} + \dots + \frac{\sigma^2}{n^2} (there are nn of these terms)
=nσ2n2= n \frac{\sigma^2}{n^2}
=σ2/n= \sigma^2 / n

Credit and licence

“A.3 Solutions: Probability” is from OpenIntro Statistics, Fourth Edition, by David M Diez, Mine Çetinkaya-Rundel and Christopher D Barr, shared under a CC BY-SA 3.0 license. WorldClass Statistics is a derivative of OpenIntro Statistics, shared under the same license; the original book may be found at openintro.org/book/os.

Source: https://www.openintro.org/book/os/

Licence: CC BY-SA 3.0 (creativecommons.org/licenses/by-sa/3.0/). You may share and adapt it under the same licence, with this credit. WorldClass adapted the formatting only: the text is the authors', with interactive parts and links to other sites left out.

Build status · Flash Infinity AI™ is a trademark of AI-Flash Inc.. WorldClass is not affiliated with, approved or accredited by any exam board, curriculum provider or university unless a page says so.